OpenStudy (shadyykatie):

1. write two equations that go through the point p(-2,5): one for a vertical line and one for a horizontal line. 2. Write an equation for the line through the point p(4, 7) with the slope of -3 in slope intercept form. 3. Find the value of x. 4. Which postulate or theorem would you use to prove each of the following pairs of triangles congruent (SSS, SAS, ASA, AAS, HL)?

OpenStudy (shadyykatie):

3:

OpenStudy (shadyykatie):

This is for #3

OpenStudy (anonymous):

1. vertical line is : x = -2 horizontal line is : y = 5 ================== 2. y = mx + b slope(m) = -3 (4,7) x = 4 and y = 7 now we sub and solve forb 7 = -3(4) + b 7 = -12 + b 7 + 12 = b 19 = b slope intercept form equation is : y = -3x + 19 =================== I am sorry but I am not good with geometry. I imagine you would use SAS for this, because you have two sides and an angle

OpenStudy (shadyykatie):

for # 4